OpenStudy (anonymous):

Expand and simplify the algebraic expression (x + 3)(x - 3) - (-x - 9)

OpenStudy (anonymous):

(x+3)(x-3)-(-x-9) =x^2+9+x+9 =x^2+x+18

jhonyy9 (jhonyy9):

x2 -9+x+9 =x2 +x =x(x+1)

OpenStudy (anonymous):

x^2+x+18

jhonyy9 (jhonyy9):

x^2 +x =x(x+1)

OpenStudy (anonymous):

i'm sory x^2+x

jhonyy9 (jhonyy9):

sure this is right sure

OpenStudy (anonymous):

same as jhonyy9

OpenStudy (anonymous):

Is what Jhonyy9 said the same as what Nuryamin said?

jhonyy9 (jhonyy9):

i am right sure --- believe me

OpenStudy (anonymous):

Oookay. Thanksss to all who answered(: